Box of doughnuts 1 2 Homers total utility 84 93 101 108 114 3 4 5 6 119 The table shows Homers total utility from boxes of
Answer #1

Marginal utility = change in Total utility

TU of 5 units = 114

TU of 6 units = 119

Marginal utility of 5th unit = 114 - 108 = 6

Marginal utility of 6th unit = 119 - 114 = 5

When consumption is reduced from 6 to 5 units Total utility decreases but marginal utility increases

option A is correct answer